OpenStudy (anonymous):

derivative of ln(xe^(7x))

OpenStudy (anonymous):

(1/(xe^(7x)))*e^(7x)*7=(7e^(7x))/(xe^(7x))=7/x

OpenStudy (anonymous):

That's not actually correct.. \[\frac{d}{dx}ln(xe^{7x}) \] \[= \frac{1}{xe^{7x}}*\frac{d}{dx}[xe^{7x}]\] \[= \frac{1}{xe^{7x}}(x\frac{d}{dx}[e^{7x}] + e^{7x}\frac{d}{dx}[x])\] \[= \frac{1}{xe^{7x}}(xe^{7x}\frac{d}{dx}[7x] + e^{7x})\] \[= \frac{1}{xe^{7x}}(7xe^{7x} + e^{7x})\] \[= \frac{7xe^{7x}}{xe^{7x}} + \frac{e^{7x}}{xe^{7x}}\] \[= 7+ \frac{1}{x}\]

OpenStudy (anonymous):

yeah, sorry, i forgot about product rule :O

OpenStudy (anonymous):

No worries =)